OpenStudy (anonymous):

What is the value of the fourth term in a geometric sequence for which a1 = 30 and r = 1/2?

OpenStudy (anonymous):

Without using the formula for the nth term you can find the 4th term easily: Geometric sequence is formed by multiplying each term by the common ratio r. 1st term = 30 2nd term = 30 * 1/2 = 15 3rd term = 15 * 1/2 = 7.5 4th term = ?

OpenStudy (anonymous):

4th term would be 7.5*1/2=3.75 ?

OpenStudy (anonymous):

thank you

OpenStudy (ilovecake):

15/4